OpenStudy (anonymous):

Solve: 6=1-2n+5 Help I don't understand any of it?

OpenStudy (anonymous):

Hi there! First of all, we need to move all of the numbers to one side and the exponents to the other. Let's subtract 6 from both sides. \[6-6=1-2n+5-6\] Now let's add 2n to both sides \[2n=1+5-6\] Make sense so far?